OpenStudy (anonymous):

17. A cone has a radius of 15 cm and a volume of 540 cm3. Find the volume of a similar cone with a radius of 10 cm. (1 point) 54 cm3 240 cm3 160 cm3 360 cm3

OpenStudy (anonymous):

V1 = 1/3 (pi) ri^2 h V2 = 1/3(pi) r2^2 h now v1/v2 = r1 ^2/ r2 ^2 540/x = (15/10)^2 = (3/2)^2 = 9/4 so x= 540*4/9 = 240

OpenStudy (anonymous):

hope u understood

OpenStudy (anonymous):

Thankk you.!

OpenStudy (anonymous):

welcome :D

OpenStudy (phi):

***a similar cone *** that means all the dimensions scale. In other words, when you shrink the radius from 15 to 10 you also shrink the height so that the smaller cone has the same shape. (if you kept the same height, the new cone would be skinnier than the original)

OpenStudy (phi):

In these problems they want you to remember that volume scales as the third power of the length of a side (or in this case radius). find the scale factor that takes you from 15 to 10 raise that same factor to the third power. multiply the original volume by this number to get the new volume. Can you do that ?

OpenStudy (phi):

find the scale factor that takes you from 15 to 10 that means find k where 15*k = 10 what is k ?